OpenStudy (anonymous):

Solve the following equation. One point is given for showing steps and one point given for the correct solution. x/x+3+2/5=0

OpenStudy (anonymous):

x = -6/7

OpenStudy (anonymous):

can u show the steps of how u got tht?

OpenStudy (anonymous):

x / x + 3 = -2/5 5x = -2(x+3) 5x = -2x - 6 7x = -6 x = -6/7

OpenStudy (anonymous):

thanks!